5. The percentage of accounting undergraduates passing the CPA exam is 20% (assume every student takes the test). You randomly select 17 graduating accounting students and would like to know a few things about their prospects of passing the test. (14 points) a) What distribution would be appropriate for this question? Why? b) What is the probability that exactly 5 will pass the exam? c) What is the probability that at least 9 will pass the exam? d) What is the probability that less than 4 will pass the exam? e) Calculate the mean, variance, and standard deviation of the number of people passing the exam in this sample?
Answer #1

a)

this is binomial distribution as events are independent and number of trails are fixed

b)

probability that exactly 5 will pass the exam =P(X=5)=17C5(0.2)5(0.8)12 =0.136076

c)

probability that at least 9 will pass the exam =P(X>=9)=0.002581

d)

probability that less than 4 will pass the exam =P(X<4)= =0.548876

e)

mean =np=17*0.2=3.4

variance =np(1-p)=2.72

std deviation =sqrt(np(1-p)=sqrt(2.72)=1.6492